You know that the price of Oriole, Inc., stock will be 24 exactly one year from today. Today the price of the stock is $21. Determine what must happen to the price of Oriole, Inc., today in order for an investor to generate a 20 percent return over the next year. Assume that Oriole does not pay dividends. to $ The price should
Answer #1

P1 = Price after a year = 24

Return to be generated after a year = 20%

So price today = P1 should be discounted by 20%

price today = P1/(1+ke) = 24/(1+0.20) = 20

price should be today 20.

So price should decrease from $21 to $20

Answer : decrease to $20 (Thumbs up please)